OpenStudy (anonymous):

Average the following birth weights: 8 Ibs. 5oz., 7lbs. 15oz., 6lbs. 3 oz., 5 lbs. 8oz., and 9lbs. 4oz.

OpenStudy (anonymous):

So what I did was: 15lbs. 20oz 20lbs 15oz. and i think its addition.

OpenStudy (anonymous):

@Jack1

OpenStudy (triciaal):

first review the other post

OpenStudy (triciaal):

try to do the same with this post so I can see your work

OpenStudy (anonymous):

ok give me a few minutes:)

OpenStudy (triciaal):

to do the average after you get the total divide by the number of entries you had to add for this problem you had 5 entries

OpenStudy (anonymous):

so this is what I got so far and got a little stuck: (15*16)+20=260+20=280 ounces (20*16)+15=335+15=350 ounces so 280+350=630 630/16=39.375 so now im stuck on where to go now

OpenStudy (triciaal):

that's what jack told you to do I prefer if you do it the easier and normal way. I am sure you will not get stuck

OpenStudy (triciaal):

continue what you had started. like in my example add all the ozs together take out groups of 16 because 16 ozs make 1 lb any remainder stays as ozs next all all the lbs including the ones you converted from groups of 16 ozs.

OpenStudy (anonymous):

im really confused how to do it the way you wrote it down. My teacher does it your way to but i just cant seem to understand it

OpenStudy (triciaal):